### v1

# The early universe is ACT-ing warm

Abstract
The recently released data from the Atacama Cosmology Telescope (ACT) confirms that the primordial scalar spectrum is extremely flat. This, together with current upper bounds on the tensor-to-scalar ratio, implies that the simplest models of inflation coming from particle physics (for instance, a minimally-coupled scalar with monomial potentials) need additional ingredients in order to make them compatible with observations. Instead of invoking arbitrary new couplings or new interactions that are not protected symmetries, we argue that dissipation of the inflaton field with the radiation bath should be added as a new physical principle. Accordingly, we show that warm inflation provides the correct paradigm to explain the current observations, given very natural choices of dissipative terms.
I Spoiled for choice, yet not a single ‘principled’ one
Our current understanding of the early universe paints a picture that is extraordinarily simple. The background spacetime is believed to have undergone an exponential expansion, commonly termed as inflation. This solves the puzzles of standard Big Bang cosmology such as why the sky looks the same in all directions or why the spatial curvature of the universe seems to be zero. However, it not only explains why the universe seems homogeneous and isotropic on large scales, but also provides the seeds for the small temperature anisotropies that we observe in the Cosmic Microwave Background (CMB) radiation (or as late-time inhomogeneities in the Baryonic Acoustic Oscillations or Large Scale Structure data).
The main prototype for inflation is a scalar field slowly rolling down an almost flat potential, such that the potential energy of the field dominates over its kinetic energy, and sources the Friedmann equation. The Hubble radius remains approximately constant during this phase, leading to an increase of the size of the universe by a factor of $e^{3N}$ if inflation lasted for $N$ $e$-folds. This results in a supercooled universe which is practically empty, and one needs a phase of reheating/preheating where the energy density of the inflaton is released by creating a shower of particles in radiation form.
The above story is simple enough that theorists have successfully built a zoo of potentials ($\sim\mathcal{O}(100)$ Martin:2013tda ) for single field models of inflation. However, if we want to go beyond pure phenomenology to construct a particle-physics motivated models of inflation, one finds that the main restriction is that the potential must remain sufficiently flat for long enough for inflation to take place. In other words, not only should the classical potential have a very flat shape, it needs to be protected from quantum corrections. This is commonly referred to as the eta-problem in inflation: How do we keep quantum radiative corrections from generating a large mass term for the inflaton? Note that the theoretical obstruction comes from noting that the potential should be sensitive to Planck-suppressed operators, the absence of which would be an extreme fine-tuning problem. However, more importantly, current observations tell us that the primordial density perturbations have an adiabatic, Gaussian and near scale-invariant spectrum. This requires that the effective field theory (EFT) of inflation is that of a light scalar field, which is precisely what the above-mentioned eta problem makes it difficult to realize.
One way to achieve the above would be protecting the flatness of the potential via symmetries, such as supersymmetry (which would cancel the bosonic and fermionic contributions to the mass) or having an approximate shift symmetry (see, for instance, Baumann:2014nda ). The first suggestion works for highly fine-tuned superpotentials McAllister:2007bg , which just kicks the naturalness problem to a different starting point. Axion fields do posses weak shift symmetries Freese:1990rb , but these are ruled out from both theoretical (super-Planckian decay constants are ruled out by the Weak Gravity Conjecture Rudelius:2015xta ; Heidenreich:2015wga ) as well as from observational constraints Stein:2021uge . For inflationary models which require a large field excursion, the sensitivity of the potential to Planckian physics is enhanced, further worsening this problem. The broader point is that since there are no global symmetries in Quantum Gravity Banks:2010zn , this means all such models employing some symmetry would require a degree of fine-tuning to make them work. A recent, systematic manner of characterizing these problems is the so-called ‘Swampland program’ Palti:2019pca ; Bedroya:2019tba .
[…]
The overarching problem of cold inflation is not that there does not exist a single model that still fits the bill as one can always reverse-engineer some potential, if required by adding more fields or additional symmetries or couplings, which can explain observed data (see Kallosh:2025rni ; Aoki:2025wld for examples post-ACT). The conceptual problem is how to screen inflation from the effects of trans-Planckian physics that come from the effects of pre-inflationary dynamics. One of the main qualities of inflation is that it is an attractor solution, and the standard Bunch-Davies initial conditions for the quantum fluctuations that source primordial inhomogeneities is a quantum attractor Kaloper:2018zgi for a quasi-de Sitter (dS) expansion. However, all of these nice qualities would be spoiled if one cannot justify how the inflaton mass remains small despite quantum corrections, as has been observed from the primordial power spectrum of density perturbations. And this requires a principled model, with a new symmetry that can be embedded from fundamental theory, which can explain why the potential remains approximately flat during inflation. Finally, note that even if such a mechanism is constructed, cold inflation would still need to invoke the reheating scenario via which the universe is repopulated giving rise to a radiation-dominated epoch at the end of inflation.
I.1 A new physical input: Dissipative corrections
The above discussion implicitly ignores the interactions of the inflaton with other matter components during the quasi-dS expansion of the universe or, at best, assumes that these interactions only affect the potential through quantum corrections. However, one should also include dissipative that typically stem from interactions of the inflaton field. If the radiation production rate is non-negligible during inflation, this would compensate the super-cooling due to the exponential expansion and would not result in an empty universe. This is what has been shown to be the case in standard models of the warm inflation (WI) paradigm Berera:1995ie , which results in a radiation bath that is in quasi-thermal equilibrium. As there exists a radiation bath, the energy of the inflaton field can be dumped into this environment continuously through dissipative terms provided the microphysics causing dissipation has a characteristic time-scale that is much larger than the local expansion rate of the background spacetime Berera:1996nv ; Berera:1998gx . As is obvious from this description, there is no need for reheating in WI as it allows a smooth and gradual phase transition from exponential expansion to a radiation-dominated standard big-bang cosmology Berera:1996fm .
Unlike in cold inflation, where the height of the potential determines the energy density of the universe while its slope fixes the slow-roll dynamics, there is an additional friction term in WI due to the radiation bath that keeps the universe warm despite the supercooling quasi-dS expansion. The background dynamics in WI is still dominated by the potential energy of the inflaton field, $V$. However, the inflaton field dynamics is not isolated and there exists the radiation energy density $\rho_{r}^{1/4}>H$, where $H$ is the approximately constant Hubble parameter during inflation. The resulting background dynamics in WI, for the inflaton and radiation energy density, modified by the presence of a dissipation term $\Upsilon$, are, respectively Berera:1995ie
$\ddot{\phi}+(3H+\Upsilon)\dot{\phi}+V_{,\phi}=0,$ (1)
$\dot{\rho}_{r}+4H\rho_{r}=\Upsilon\dot{\phi}^{2},$ (2)
where $\Upsilon$ is the dissipation coefficient (for reviews on its derivation from first principles quantum field theory, see, e.g. Berera:2008ar ; Kamali:2023lzq and references there in) and it can be, in general, a function of both the temperature $T$ of the generated thermal radiation bath and on the inflaton amplitude $\phi$.
For WI to take place, the energy density of the inflaton field must dominate over that of the radiation field. On introducing the dissipation coefficient $Q=\Upsilon/(3H)$, one can rewrite standard slow-roll parameters as
$\epsilon_{V}$ $=$ $\frac{M_{\rm Pl}^{2}}{2}\left(\frac{V^{\prime}}{V}\right)^{2}<1+Q\,,$ (3)
$\eta_{V}$ $=$ $M_{\rm Pl}^{2}\left(\frac{V^{\prime\prime}}{V}\right)<1+Q\,$ (4)
where $M_{\rm Pl}$ is the reduced Planck mass and $Q$ is the dissipation ratio $Q=\Upsilon/(3H)$. It is clear that if WI allows for strong dissipation, $Q\gg 1$, then the problem of requiring that the inflaton mass remains negligible during inflation is automatically solved. Put another way, the strong dissipation allows for a slow-roll trajectory by proving additional friction even if quantum corrections somewhat spoil the flatness of the potential.
The presence of the dissipation term in WI not only modifies the background dynamics, but also strongly affects the perturbations, with in particular the inflaton and radiation perturbations becoming coupled due to the temperature dependence on $\Upsilon$ Berera:1995wh ; Berera:1999ws . In this case, the scalar of curvature of perturbations in WI is modified with respect to the CI case and it can be written as Ramos:2013nsa ; Benetti:2016jhf
$P_{\mathcal{R}}(k/k_{*})=\left(\frac{H^{2}}{2\pi\dot{\phi}}\right)^{2}\mathcal{F}(k/k_{*}),$ (5)
where all quantities are meant to be evaluated at Hubble radius crossing scale $k=aH$. In Eq. (5), $\mathcal{F}(k/k_{*})$ is defined as
$\mathcal{F}(k/k_{*})\equiv\left(1+2n_{*}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)G(Q),$
where $n_{*}$ accounts for the possibility of a thermal distribution of the inflaton field due to the presence of the radiation bath and assumes the Bose-Einstein distribution $n_{*}\equiv n_{\rm BE}=1/[\exp(H/T)-1]$ when fully thermalized (for intermediate cases, see, e.g. Ref. Bastero-Gil:2017yzb ). In Eq. (6), the function $G(Q)$ models the effect of the coupling of the inflaton and radiation fluctuations and which can only be determined numerically by solving the coupled system of perturbation equations in WI Graham:2009bf ; Bastero-Gil:2011rva ; Bastero-Gil:2014jsa . Equation (5) can be expanded in terms of the small scale $k$-dependence as
$P_{\mathcal{R}}(k/k_{*})\simeq A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1+\frac{\alpha_{s}}{2}\ln(k/k_{*})+\cdots},$ (7)
where $A_{s}$ is the scalar amplitude, $n_{s}$ is the scalar tilt,
$n_{s}-1=\frac{d\ln P_{{\cal R}}(k/k_{*})}{d\ln(k/k_{*})}\Bigr{|}_{k\to k_{*}},$ (8)
and $\alpha_{s}$ is the running (the ellipses in Eq. (7) denote higher order running terms, which we neglect here),
$\alpha_{s}=\frac{dn_{s}(k/k_{*})}{d\ln(k/k_{*})}\Bigr{|}_{k\to k_{*}}.$ (9)
In WI, expressions for the tilt and running can be obtained explicitly by using Eq. (5) (see also, e.g.  Das:2022ubr ).
The Planck+ACT+DESI data (P-ACT-LB hereafter) gives the best fit value of $n_{s}=0.974\pm 0.003$, which clearly shows that the spectrum is even more scale invariant. If one adds in the BICEP/Keck (P-ACT-LB-BK18 hereafter) data BICEP:2021xfz ; BICEPKeck:2024stm , one can turn this into a constraint on the second slow-roll parameter $\eta_{V}\sim\mathcal{O}(10^{-3})$ ACT:2025fju ; ACT:2025tim , which is very close to zero. Thus, the almost scale-invariance of the spectrum can be seen to make the eta-problem of cold inflation even worse, requiring extreme fine-tuning of the alpha-attractor models to include further arbitrary parameters of non-minimal couplings of the inflaton to gravity Kallosh:2025rni ; Kallosh:2013tua . On the other hand, this is not required in WI. For the strong dissipation regime, the value of $n_{s}$ is naturally led to larger values closer to one. The physical reason behind this is that a given wavelength mode attains its constant value slightly before it exits the horizon while the dissipative terms slows down the evolution of the field.
Although quantum fluctuations are omnipresent in all models of inflation, Eq. (5) shows that the thermal and dissipation effects can dominate over them in WI and are primarily responsible for sourcing the adiabatic curvature perturbations. More interestingly, in the strong dissipative regime, where $Q>1$, these fluctuations freeze out even before they become superhorizon leading to an enhancement in the amplitude of the power spectrum. Since there is no such modification to the spectrum of the primordial tensor modes, this means that WI models would always lead to a suppression of the tensor-to-scalar ratio $r$ (noticing that dissipative effects to the gravitons are tiny for any temperature below the Planck scale). This implicitly is another reason why data seems to be favouring WI models, as the upper bound on $r$ is now constrained to be below $0.036$ BICEP:2021xfz . This will be further constrained in the near future by various missions such as LiteBird, BICEP/Keck, CMB-S4 etc. How large the value of $r$ we find would be the smoking gun in deciding between weak and strong dissipative regimes of WI, and would be crucial as an indication of whether the ‘natural’ models of WI, parametrized by a large value of $Q$, are preferred or not.
II It is time to ACT now
P-ACT-BL has solidified the trend of pushing $n_{s}$ to higher values while combining the BK-18 results on a tiny upper bound on the tensor-to-scalar ratio. We would have to wait until the next data release of South Pole Telescope (SPT) to see if this trend of increasing $n_{s}$ while other missions will further constrain the upper bound on the amplitude of primordial gravitational waves. Our main message is that one does need to keep on fine-tuning existing cold inflation models by further adding parameters to account for future observations. Instead a principled approach would be to see what would happen if we were to add the dissipative corrections coming from WI and see what they predict the typical values for $n_{s}$ and $r$ are.
The warm inflation model building program has focused on models built with the types of fields and symmetries that are as close as possible to those already found in the Standard Model, and have thus been experimentally observed. Developing such a goal has been possible due to two key steps. The first, we made in the early years of development of WI was the recognition that when the dissipative coefficient $\Upsilon$ is larger than the Hubble scale $H$, i.e. $\Upsilon>3H$, it is possible in this strong dissipative regime for the inflaton mass to be larger than the Hubble scale, $m_{\phi}>H$, and the inflaton field amplitude to remain below the Planck scale $\phi<m_{pl}$ Berera:1999ws ; Berera:2003yyp . Such a regime is then protected from quantum gravity corrections arising in higher dimensional operators and free of uncontrollable infrared problems. This is the minimal regime necessary to be consistent with the quantum field theory understood by the Standard Model. The second key step was recognizing, after much research by us, that for WI it is not Supersymmetry that best protects the inflaton mass from radiative and thermal corrections but rather the symmetries associated with the Nambu-Goldstone boson, leading to the first WI model based on this symmetry, the Warm Little inflaton model Bastero-Gil:2016qru . These two steps created a new path for inflation model building and the one followed in most WI model building today (for a more detailed discussion please see Berera:2023liv ).
II.1 The models
The first model we consider is the warm little inflaton model introduced in Ref. Bastero-Gil:2016qru , with the inflaton having direct interactions with fermion fields (we are going to refer to this model as WLIF for short) and with a quartic inflaton potential and interaction Lagrangian density given by
$\mathcal{L}_{\rm int}$ $=$ $-{g\over\sqrt{2}}(\phi_{1}+\phi_{2})\bar{\psi}_{1L}\psi_{1R}+i{g\over\sqrt{2}}(\phi_{1}-\phi_{2})\bar{\psi}_{2L}\psi_{2R}$ (10)
-- $h\sigma\sum_{i=1,2}\left(\bar{\psi}_{iL}\psi_{\sigma R}+\bar{\psi}_{\sigma L}\psi_{iR}\right),$
$\phi_{1}$ and $\phi_{2}$ are two complex scalar fields, with vacuum expectation values, $\langle\phi_{1}\rangle=\langle\phi_{2}\rangle\equiv M/\sqrt{2}$ and that are charged under the same $U(1)$ symmetry. The common phase is a pseudo-Goldstone boson that can be taken as the inflaton field. The remaining field content in Eq. (10), $\sigma$ is a scalar singlet, $\psi_{\sigma R}$ are chiral fermions carrying the same charge of $\psi_{1}$ and $\psi_{2}$, and $\psi_{\sigma L}$ has zero charge. The model satisfies the symmetries $\phi_{1}\leftrightarrow i\phi_{2}$ and $\psi_{1L,R}\leftrightarrow\psi_{2L,R}$, which protects the inflaton potential to receive any large quantum or thermal corrections. The dissipation coefficient evaluated in this case is
$\Upsilon=C_{T}T~{},\qquad C_{T}\simeq\alpha(h)g^{2}/h^{2}~{},$ (11)
where $\alpha(h)\simeq 3/[1-0.34\ln(h)]$. In the WLIF model, the form of the dissipation coefficient leads to a growing function $G(Q)$ in the power spectrum, which in general drives the tilt faster to a blue value ($n_{s}>1$). The simplest inflaton potential that can be considered in this case is a quartic monomial potential,
$V(\phi)=\frac{\lambda}{4}\phi^{4},$ (12)
which we will consider when presenting the results for this model.
The second model we analyze is a variant version of the WLIF model, with interaction Lagrangian density given by Bastero-Gil:2019gao
$\mathcal{L}_{\rm int}$ $=$ $-{1\over 2}g^{2}|\phi_{1}+\phi_{2}|^{2}|\chi_{1}|^{2}-{1\over 2}g^{2}|\phi_{1}-\phi_{2}|^{2}|\chi_{2}|^{2}$
$+$ $\!\!\!\!\!\sum_{\begin{subarray}{c}i\neq j=1,2\end{subarray}}\!\!\!\left(h\chi_{i}\bar{\psi}_{L}\psi_{R}\!+\!\mathrm{h.c.}\!-\!{\lambda\over 2}|\chi_{i}|^{4}\!-\!{\lambda^{\prime}}|\chi_{i}|^{2}|\chi_{j}|^{2}\!\right)\!,$
with principal scalar interactions instead of the femrion ones of the first model. We will denote this model by WLIS for short. As in the first model, the symmetries prevent the inflaton potential again from receiving large corrections. The dissipation coefficient derived in this case is found to be given by In this scalar field variant of the WLI model, the dissipation coefficient now becomes Bastero-Gil:2019gao
$\Upsilon\simeq{4g^{4}\over h^{2}}{M^{2}T^{2}\over\tilde{m}_{\chi}^{3}}\left[1+{1\over\sqrt{2\pi}}\left({\tilde{m}_{\chi}\over T}\right)^{3/2}\right]e^{-\tilde{m}_{\chi}/T},$ (14)
where $\tilde{m}_{\chi}^{2}\simeq g^{2}M^{2}/2+\alpha^{2}T^{2}$, with $\alpha^{2}\simeq\left[h^{2}+2\lambda+\lambda^{\prime}\right]/12$. In the WLIS model, the dissipation coefficient behaves with a negative power in the temperature, $\Upsilon\propto 1/T$ for appropriate choices of parameters of the model. In this case, the function $G(Q)$ behaves like a decreasing function with the increase of $Q$. The simplest potential that can be studied in this case is just a quadratic monomial potential,
$V(\phi)=\frac{m_{\phi}^{2}}{2}\phi^{2},$ (15)
which is the choice of potential we consider here for the WLIS model.
In the next section we present our results for these two models and contrast them with the recent P-ACT-LB-BK18 data.
II.2 Results
All of the results we are going to present are obtained by solving the WI equations given in the previous section, with the appropriate forms for the function $G(Q)$ correcting the power spectrum Eq. (6) which are derived for each of models considered (see, e.g. Refs. Graham:2009bf ; Bastero-Gil:2011rva for details and also Ref. Montefalcone:2023pvh for a public code to numerically determine $G(Q)$ for different forms of dissipation coefficients in WI). We also study the two limiting cases of considering the inflaton perturbations with negligible thermalization or fully thermalized, taking $n_{*}=0$ or $n_{*}=n_{BE}$ in Eq. (6). Finally, by setting the pivot scale, the point (given by the number of e-folds $N_{*}$) during inflation where it crosses the Hubble radius is determined explicitly by Das:2020xmh
$\frac{k_{*}}{a_{0}H_{0}}=e^{-N_{*}}\left[\frac{43}{11g_{s}(T_{\rm end})}\right]^{1/3}\frac{T_{0}}{T_{\rm end}}\frac{H_{*}}{H_{0}}\frac{a_{\rm end}}{a_{\rm reh}},$ (16)
[…]
Figure 1: The one-sigma and two-sigma constrains from Planck (orange shaded regions) and the combined P-ACT-LB results (violet shaded region) for $r$ and $n_{s}$, together with the results for the WLIF (panel a) and WLIS (panel b) models. The curves are the results obtained at different values for the dissipation ratio $Q$ at Hubble crossing (marked by the dots shown in the figure). The parameters considered in each case are the same as those used in Bastero-Gil:2016qru (for the WLIF model) and Bastero-Gil:2019gao (for the WLIS model).
In Fig. 1, we show the results for the two models, the WLIF and WLIS, obtained for the spectral tilt $n_{s}$ and for the tensor-to-scalar ratio $r$ as the value of $Q_{*}$ (evaluated at $N_{*}$) is varied (which corresponds to changes in the parameters and couplings of the two models). The behavior seen in the curves for the WI models in the figures in $r\times n_{s}$ is a consequence of the relative importance of each term in the power spectrum Eq. (full-power and how $Q$ and $T/H$ evolve (see, e.g. Ref. Das:2020lut for a discussion about the dynamical evolution for $Q$ and $T/H$ in WI and for their effect on $n_{s}$, see Ref. Das:2022ubr ).
[…]
Figure 2: Similar to the Fig. 1, but now showing the running $\alpha_{s}$ as a function of the dissipation ratio $Q_{*}$ for the WLIF and WLIS models.
In addition to finding a slight spectral tilt towards scale invariance compared to Planck combined with BICEP/Keck data BICEP:2021xfz , the recent ACT data tend to slight favor a positive running (though still compatible with a null value), $\alpha_{s}=0.0062\pm 0.0052$. It is interesting to see that WI can also easily accommodate such values. The physical reason behind this is that the negative values for $\alpha_{s}$ in cold models come arise from the fact that the inflationary trajectories deviate from quasi-dS at the end of inflation when the shortest wavelength modes exit the horizon. While for WI, the energy density of the inflaton decreases continuously leading to a smooth transition to radiation domination. In Fig. 2, we show the results for the two models, the WLIF and WLIS, for the running $\alpha_{s}$ in terms of spectral tilt $n_{s}$, once again, different points in the curves correspond to different values of $Q_{*}$. In each case, the solid line corresponds to the results when neglecting $n_{*}$ in the equation for the scalar spectrum in WI or taking $n_{*}$ as a Bose-Einstein distribution (dashed line). The results fall well inside the one-sigma region found by ACT ACT:2025tim . It also show that for large $Q$ (as $n_{s}$ increases), the running tends to move from negative (in the weak dissipation regime of WI) towards positive values (corresponding to the strong dissipative WI regime), but still remains very small.
III Conclusions
Where does the P-ACT-LB-BK18 data leave inflationary model-building? At face value, a tiny $r$ and $n_{s}$ very close to unity tells us that the standard models of renormalizable, monomial potentials seem to fall out of favour of observations. However, the deeper revelation is that observations seem to strike at the heart of the problem of protecting the flatness of the potential from quantum corrections. One can always find one or two-parameter extensions of current models of cold inflation which fit the data as more precise mappings of the CMB come in; however, it is time we introspect to ask if such extensions have an embedding in any UV-theory. Put another way, can one accentuate well-motivated models with such ad hoc couplings which are neither derived from any underlying theory nor protected from the effect of Planck-suppressed operators? Even when one takes the modern view of EFT that non-renormalizable potentials are, by themselves, not problematic and simply point towards a UV-cutoff to which scale it can be trusted, it makes little sense to pick and choose operators (of the same derivative order) such that a model becomes viable. A principled path forward for inflation seems to require a new physical input which would help reconciling theoretically-consistent models with observations.
In this paper, we argue that WI is the right paradigm to do so. It is only natural that the inflaton field interacts with other fields during inflation and the only further assumption WI makes is regarding the strength of the dissipative terms generated in the process. However, one then finds that standard models of WI can easily fit the data where their cold counterparts failed. One might question why is the specific potentials chosen in this paper not equally fine-tuned. Our main argument is that the exact choice of potential is not as important in WI as it is in the cold case. This is precisely because of the presence of the extra dissipative terms. We could just as easily have chosen some other potential and much of our discussion would still go through. Moreover, we do choose a form of the potential which is renormalizable and thereby can be assumed to be protected from Planckian corrections. As can be clearly seen from Fig. 1, dissipative corrections turn even these monomial potentials viable in WI. Notice that although the strong dissipative regime would be preferred from theoretical considerations, these models produce the right spectrum even for small $Q$. However, future observations of the tensor-to-scalar ratio, or the lack of it, can be used as a signature of the strong vs. weak dissipative regime for WI.
Naturally, WI is also not free from any constraints. Firstly, the dissipative coefficients need to be computed from first-principles (and that is what has been done for the two models under consideration here). More needs to be done regarding the microphysics of such dissipative interactions and how they can be derived from a full thermal field theory perspective in an accelerating background. This can put further bounds on the parameter space of WI. Moreover, $Q$ is not a completely unconstrained parameter even from observations. It is natural to expect that going to arbitrarily large values of the dissipative coefficient would lead to an unacceptably large $f_{NL}$ for the warm shape in the scalar bispectrum (we have made sure that going upto $Q\sim\mathcal{O}(100)$ is consistent with current bounds on the bispectrum Bastero-Gil:2019gao ). We plan to find the upper bound on $Q$ coming from the bispectrum, which would also restrict how large $n_{s}$ can be in WI.
For completeness, we also mention the possibility that inflation itself might not be the right way to think about the early-universe. A few alternatives to inflation Brandenberger:1988aj ; Steinhardt:2001st ; Agrawal:2020xek ; Brahma:2022dsd ; Boyle:2022lcq , which are grounded in a fundamental UV-theory, might be better suited to explain observations. Curiously, many of them also postulate thermal fluctuations Nayeri:2005ck ; Brahma:2021tkh as in WI, thereby pointing towards a larger application of dissipative corrections going beyond inflation.
Acknowledgements.
AB is partially funded by STFC. SB is supported in part by the Higgs Fellowship and by the STFC Consolidated Grant “Particle Physics at the Higgs Centre”. R.O.R. and G.S.R. acknowledge financial support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) - Finance Code 001. R.O.R. also acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.